the perimeter of a rectangle is 220 inches. The length exceeds the width by 78 inches. find the length and width.
The length is______inches.
(Type and integer or a decimal)

Answer by phxladyphx@yahoo.com(18)

P=perimeter
x=width
x+78=length
we know the perimeter is 220 inches and we know that L+W+L+W=P
x+78+x+78+x+x=220
4x+156=220
4x+156-156=220-156
4x=64
4x/4=64/4
x=16
Width is x so the width is 16 inches.
Length is width plus 78 inches.
16+78=94
The length is 94 inches.